### v2

# Lizard: An Efficient Linearization Framework for Large Language Models

Abstract
We propose Lizard, a linearization framework that transforms pretrained Transformer-based Large Language Models (LLMs) into flexible, subquadratic architectures for infinite-context generation. Transformer-based LLMs face significant memory and computational bottlenecks as context lengths increase, due to the quadratic complexity of softmax attention and the growing key-value (KV) cache. Lizard addresses these limitations by introducing a subquadratic attention mechanism that closely approximates softmax attention while preserving the output quality. Unlike previous linearization methods, which are often limited by fixed model structures and therefore exclude gating mechanisms, Lizard incorporates a gating module inspired by recent state-of-the-art linear models. This enables adaptive memory control, supports constant-memory inference, offers strong length generalization, and allows more flexible model design. Lizard combines gated linear attention for global context compression with sliding window attention enhanced by meta memory, forming a hybrid mechanism that captures both long-range dependencies and fine-grained local interactions. Moreover, we introduce a hardware-aware algorithm that accelerates the training speed of our models. Extensive experiments show that Lizard achieves near-lossless recovery of the teacher model’s performance across standard language modeling tasks, while significantly outperforming previous linearization methods. On the 5-shot MMLU benchmark, Lizard improves over prior models by 18 points and shows significant improvements on associative recall tasks.
1 Introduction
Large Language Models (LLMs) based on Transformer architectures have achieved impressive advances in natural language processing across various tasks (Grattafiori et al., 2024; Mann et al., 2020; Achiam et al., 2023). However, their reliance on softmax attention with quadratic time and memory complexity poses significant limitations for long-context applications, both during training and inference. In particular, the softmax attention computation scales quadratically with the sequence length, while the key-value (KV) cache grows linearly during generation, resulting in significant computational and memory overhead for long-context sequences (Adnan et al., 2024).
Recent work has proposed linear and subquadratic alternatives to softmax attention (Yang et al., 2023; Gu and Dao, 2023; Peng et al., 2023b; Wang et al., 2020; Dao and Gu, 2024), enabling linear-time training and constant-memory inference. However, despite these efficiency benefits, pretraining LLMs with such architectures from scratch requires massive training budgets, often involving trillions of tokens. More importantly, models trained with linear attention mechanisms from scratch consistently underperform on tasks that require in-context learning and retrieval. For example, Transformer models significantly outperform both Mamba and Mamba-2 by up to 15 points on the 5-shot MMLU benchmark when all are pretrained under the same settings (Waleffe et al., ). Another promising direction involves linearizing pretrained Transformer-based LLMs by replacing their softmax attention modules with subquadratic alternatives (Zhang et al., ; Mercat et al., 2024; Zhang et al., 2024; Wang et al., 2024). Instead of training new models from scratch with linear attention, this approach leverages the capabilities of openly available, large-scale pretrained models typically trained on trillions of tokens, and retrofits them with more efficient attention mechanisms. By starting from high-quality pretrained models, linearized models can better preserve task performance and recover softmax-like spiky attention patterns, thereby enhancing retrieval performance in in-context learning tasks (Zhang et al., 2024).
However, prior work either fails to fully exploit the strengths of linear attention mechanisms or still significantly underperforms compared to softmax-based attention. For example, LoLCATs (Zhang et al., 2024), a recent state-of-the-art linearization method, shows promising results and cost-effectiveness. However, it still relies on predefined positional embedding patterns for positional encoding, thereby struggling with length generalization and failing to extrapolate beyond the trained context length. Moreover, due to architectural constraints and limited flexibility, it lacks an efficient memory update mechanism, which significantly affects its performance on recall-intensive tasks. On the other hand, models such as MambaInLlama (Wang et al., 2024) show better extrapolation to longer sequences, but still suffer from substantial performance degradation. Specifically, on the 5-shot MMLU benchmark, LoLCATs lags behind its teacher model by 13.8 points, while MambaInLLaMA exhibits a 23.4 point drop.
In this paper, we introduce Lizard (Linearizing Softmax Attention with Recurrent Gate Dynamics), an efficient framework for linearizing LLMs. Unlike prior approaches that constrain themselves to reusing only the pretrained model’s existing parameters and fixed architecture (Mercat et al., 2024; Zhang et al., 2024), we introduce additional learnable components in the form of a gating module. The gating module serves two primary purposes. First, it acts as a data-dependent alternative to predefined positional embedding methods, such as Rotary Positional Embeddings (RoPE) (Heo et al., 2024), allowing the model to learn implicit relative positional information through adaptive decay patterns, enhancing length generalization. Second, the gated recurrent structure provides a dynamic memory management mechanism, enabling the model to control the retention and forgetting of past tokens during generation, thereby improving associative recall.
To further recover the expressiveness of softmax attention, we combine the output of the gated linear attention with a sliding window attention mechanism augmented with meta memory. While the gated component captures global context in a compressed form, the sliding window attention preserves local precision and expressivity within the local context. This hybrid attention framework provides a high-quality approximation to softmax attention. We train our model in two stages: (1) an attention approximation stage where the subquadratic modules are trained to mimic softmax attention outputs, and (2) a fine-tuning stage where the linearized model is adapted to downstream language modeling objectives. Our contributions can be summarized as follows:
  • We propose Lizard, a linearization framework that adapts pretrained Transformer-based LLMs into gated recurrent architectures. This adaptation enables a highly efficient, subquadratic attention mechanism with constant-memory inference and strong length generalization. Lizard achieves near-lossless recovery of the original model’s performance across major benchmarks and outperforms prior linearization methods by 8–18 points on the 5-shot MMLU benchmark.
  • We conduct extensive empirical studies to analyze architectural design choices within Lizard and evaluate which state-of-the-art recurrent architectures are most effective, across diverse benchmarks.
  • We present a hardware-aware implementation that improves the training throughput by up to 24%, enabling faster optimization.
2 Preliminary
To motivate Lizard, we first review the core components of Causal Softmax Attention, Gated Linear Attention, and techniques for approximating softmax attention. We then briefly discuss related work on linearizing Transformers and adapting pretrained LLMs for efficient long-context modeling.
Causal Softmax Attention In modern Transformer architectures (Touvron et al., 2023; Jiang et al., ), for a query vector at position $i$, Causal Softmax Attention with RoPE produces the output $\mathbf{y}_{i}$ as:
$\mathbf{y}_{i}=\sum_{t=1}^{i}\frac{\exp\left(\varphi_{\text{RoPE}}(\mathbf{q}_{i})^{\top}\varphi_{\text{RoPE}}(\mathbf{k}_{t})/\sqrt{d}\right)}{\sum_{j=1}^{i}\exp\left(\varphi_{\text{RoPE}}(\mathbf{q}_{i})^{\top}\varphi_{\text{RoPE}}(\mathbf{k}_{j})/\sqrt{d}\right)}\mathbf{v}_{t},$
where $\varphi_{\text{RoPE}}(\mathbf{q}_{i})$ and $\varphi_{\text{RoPE}}(\mathbf{k}_{i})$ denote RoPE-transformed query and key vectors, respectively.
The causal softmax attention with RoPE can be expressed in matrix multiplication form as below:
$\mathbf{Y}_{\text{softmax}}=\mathrm{softmax}\left(\mathbf{Q}_{\text{RoPE}}\mathbf{K}_{\text{RoPE}}^{\top}/{\sqrt{d}\odot\mathbf{M}}\right)\mathbf{V}$
where $L$ is the context length, $\mathbf{Q}_{\text{RoPE}},\mathbf{K}_{\text{RoPE}}\in\mathbb{R}^{L\times d}$ are the RoPE-transformed query and key matrices, $M\in\mathbb{R}^{L\times L}$ is the lower-triangular causal mask ($\mathbf{M}_{ij}=1$ if $i\geq j$, and $\mathbf{M}_{ij}=-\infty$ if $i<j$), $\mathbf{V}\in\mathbb{R}^{L\times d}$ is the value matrix, $\odot$ denotes element-wise multiplication.
Gated Linear Attention Linear attention mechanisms (Katharopoulos et al., 2020) approximate softmax attention by replacing the exponential similarity function with a kernel function $k(q,k)=\phi(q)\phi(k)^{\top}$, where $\phi$ is a feature map. Gated Linear Attention (GLA), introduced by (Yang et al., 2023), is a variant of linear attention that incorporates a learnable decay term. The output of gated linear attention $\hat{\mathbf{y}}_{i}$ is computed as:
$\hat{\mathbf{y}}_{i}=\frac{\phi_{q}(\mathbf{q}_{i})^{\top}\sum_{t=1}^{i}\left(\prod_{l=t+1}^{i}\Gamma_{l}\right)\phi_{k}(\mathbf{k}_{t})\mathbf{v}_{t}}{\phi_{q}(\mathbf{q}_{i})^{\top}\sum_{j=1}^{i}\left(\prod_{l=j+1}^{i}\Gamma_{l}\right)\phi_{k}(\mathbf{k}_{j})},$
where $\Gamma_{i}=\text{sigmoid}(W_{\gamma}\mathbf{x}_{i})$ is a data-dependent gating factor with $W_{\gamma}\in\mathbb{R}^{d\times d}$. The term $\prod_{l=t+1}^{i}\Gamma_{l}$ represents a decaying factor applied to the contribution from position $t$. Critically, this formulation enables a recurrent structure that allows constant memory inference:
$\mathbf{S}_{i}=\Gamma_{t}\mathbf{S}_{i-1}+\phi_{k}(\mathbf{k}_{i})\mathbf{v}_{i}^{\top},\quad\hat{\mathbf{y}}_{i}=\mathbf{S}_{i}\phi_{q}(\mathbf{q}_{i})$
Let $\mathbf{c}_{t}=\prod_{j=1}^{t}\Gamma_{t}$ and define $\mathbf{C}\in\mathbb{R}^{L\times d}$ as the matrix obtained by stacking the vectors $\mathbf{c}_{t}$. The GLA output can then be expressed in matrix form as:
$\hat{\mathbf{Y}}_{gate}=\left(\left(\left(\phi_{q}(\mathbf{Q})\odot\mathbf{C}\right)\,\left(\frac{\phi_{k}(\mathbf{K})}{\mathbf{C}}\right)^{\top}\right)\odot\mathbf{M}\right)\mathbf{V}$ (1)
Linearizing Transformers To combine efficiency with quality, (Zhang et al., ) proposes a learnable feature map that preserves the spiky and monotonic properties of softmax attention, crucial for selective focus, while maintaining linear time complexity. The feature map Hedgehog is defined as:
$\phi(\mathbf{x})=\left[\exp(\mathbf{x}W)\oplus\exp(-\mathbf{x}W)\right],$
where $\oplus$ is the concatenation operation. Hedgehog demonstrated that the symmetric exponential form enables low-entropy, high-selectivity attention patterns. Prior methods have exploited Hedgehog as a feature map to linearize Transformer-based LLMs (Zhang et al., 2024).
3 Lizard: Subquadratic Attention for Infinite Context
Figure 1: Overview of the Lizard training pipeline. The model is trained in two stages. Stage 1: The pretrained model is frozen, and the added modules are trained to approximate both the positional encoding patterns and the softmax attention outputs. Stage 2: The softmax attention is replaced with the trained hybrid modules, then fine-tuned on the next-word prediction objective.
To enable infinite-context generation, we replace softmax attention with a subquadratic mechanism by introducing learnable modules that approximate both attention weights and RoPE patterns. This enables constant-memory inference while preserving generalization to longer sequences.
Prior linearization approaches typically retain the original architecture of pretrained models and avoid adding new parameters (Mercat et al., 2024; Zhang et al., 2024), which limits their flexibility and prevents the use of more expressive designs. As a result, they inherit architectural constraints that hinder performance and generalization. In contrast, we introduce a learnable gating module inspired by recent advances in gated recurrent architectures (Yang et al., 2023). This design choice allows for adaptive memory control, stronger length generalization, and more effective recovery of the original model’s capabilities.
To recover the performance of softmax attention, we adopt a two-stage training procedure (Figure 1). In the First Stage, we remove RoPE and replace softmax attention with a hybrid subquadratic mechanism. The added modules are trained to mimic softmax behavior, including both attention weights and RoPE-induced positional patterns, enabling generalization to longer contexts. In the Second Stage, we replace the original attention layers with the trained subquadratic modules and employ fine-tuning to align the them with downstream language modeling objectives.
3.1 First Stage: Approximating Softmax Attention for Unbounded Context
In the first stage, we introduce three new components: two Hedgehog feature maps, $\phi_{q}$ and $\phi_{k}$, and a learnable gating module parameterized by $W_{\gamma}$. These are integrated into the original attention mechanism to form Gated Linear Attention (GLA), which serves as a subquadratic approximation of the original softmax attention. We approximate the causal softmax attention with RoPE with GLA:
$\mathbf{y}_{i}=\sum_{t=1}^{i}\frac{\exp\left(\varphi_{\text{RoPE}}(\mathbf{q}_{i})^{\top}\varphi_{\text{RoPE}}(\mathbf{k}_{t})/\sqrt{d}\right)\mathbf{v}_{t}}{\sum_{j=1}^{i}\exp\left(\varphi_{\text{RoPE}}(\mathbf{q}_{i})^{\top}\varphi_{\text{RoPE}}(\mathbf{k}_{j})/\sqrt{d}\right)},\quad\hat{\mathbf{y}}_{i}=\frac{\phi_{q}(\mathbf{q}_{i})^{\top}\sum_{t=1}^{i}\left(\prod_{l=t+1}^{i}\Gamma_{l}\right)\phi_{k}(\mathbf{k}_{t})\mathbf{v}_{t}}{\phi_{q}(\mathbf{q}_{i})^{\top}\sum_{j=1}^{i}\left(\prod_{l=j+1}^{i}\Gamma_{l}\right)\phi_{k}(\mathbf{k}_{j})}$
where $\Gamma_{i}=\text{sigmoid}(W_{\gamma}\mathbf{x}_{i})$ is a data-dependent gating factor with $W_{\gamma}\in\mathbb{R}^{d\times d}$
The gating mechanism plays a dual role in the attention transformation. First, it implicitly captures relative positional information by controlling the decay of past contributions. Unlike RoPE, which relies on predefined sinusoidal patterns, the data-adaptive gating factors $\gamma_{i}$ enable better generalization across context lengths. Second, the gating mechanism supports a recurrent formulation that allows constant-memory inference. The accumulated hidden state $\mathbf{S}_{i}$ incrementally summarizes historical information up to position $i$ using gated updates, removing the need to store the full key-value sequence.
Hardware-aware implementation The matrix form of gated linear attention in Eq. 1 is not numerically stable, as the cumulative product of gating values $\mathbf{c}_{t}=\prod_{j=1}^{t}\Gamma_{j}$ can become extremely small, leading to underflow and instability during training. Prior work mitigates this by operating in log-space and introducing a secondary-level chunking mechanism to regain compatibility with tensor core matrix multiplications (Yang et al., 2023). However, this method still requires full-precision computation within chunks, limiting overall efficiency.
To address this, we propose a more hardware-efficient factorization that directly incorporates the log-gating values into the query and key projections. Specifically, we reparameterize the attention computation by absorbing the cumulative gates into the queries and keys as:
$\widetilde{Q}$ $=\phi_{q}(\mathbf{Q})\odot\mathbf{C}=\phi_{q}(\mathbf{Q})\odot\exp(\log\mathbf{C})$
$\widetilde{K}$ $=\frac{\phi_{k}(\mathbf{K})}{\mathbf{C}}=\phi_{k}(\mathbf{K})\odot\exp(\log-\mathbf{C})$
Substituting these into Eq. 1, the gated attention becomes: $\mathbf{Y}_{\text{gate}}=\left(\left(\widetilde{Q}\,\widetilde{K}^{\top}\right)\odot\mathbf{M}\right)\mathbf{V}.$
This reparameterization is only numerically stable when the feature maps $\phi_{q}$ and $\phi_{k}$ are nonnegative. In our case, we leverage the non-negativity of the Hedgehog feature map, enabling us to express the gated projections compactly as:
$\widetilde{Q}$ $=\left[\exp(\mathbf{Q}W+\log\mathbf{C})\oplus\exp(-\mathbf{Q}W+\log\mathbf{C})\right],$
$\widetilde{K}$ $=\left[\exp(\mathbf{K}W-\log\mathbf{C})\oplus\exp(-\mathbf{K}W-\log\mathbf{C})\right].$
By shifting the gating contributions into the feature space, our formulation enables stable and parallelizable matrix operations compatible with tensor core units. This allows for an efficient chunkwise GPU implementation with no need for full-precision fallback. Empirically, our method yields stable training dynamics and achieves up to a 32% speedup compared to prior gated attention implementations [^1] https://github.com/fla-org/flash-linear-attention (Yang et al., 2023).
Figure 2: Visualize attention map of Lizard, as a combination of Gated Linear Attention and Sliding Window Attention with Meta Memory (with $w=3$ and $m=2$)
Sliding Window Attention with Meta Memory Inspired by prior work demonstrating that hybrid combinations of attention layers and linear attention mechanisms can improve model quality (Arora et al., 2024; Munkhdalai et al., 2024), we incorporate a short sliding window attention (SWA) module combined with a meta-memory mechanism. While gated linear attention captures compressed global context, SWA preserves the expressive nonlinearity of softmax attention, enabling effective modeling of fine-grained local dependencies. Specifically, for a window size $w$, each query vector $\mathbf{q}_{i}$ attends only to the most recent $w$ keys $\mathbf{k}_{i-w+1},\dots,\mathbf{k}_{i}$}, along with a small set of $m$ prepended special "meta" tokens $\mathbf{k}_{0},\dots,\mathbf{k}_{m-1}$, which serve as compressed global context representations. Formally, the output of the sliding window attention at position $i$ is computed as:
$\hat{y}_{i}=\frac{\sum_{t\in\{0,\dots,m-1\}\cup\{i-w+1,\dots,i\}}\exp(\mathbf{q}_{i}^{\top}\mathbf{k}_{t}/\sqrt{d})\mathbf{v}_{t}}{\sum_{t\in\{0,\dots,m-1\}\cup\{i-w+1,\dots,i\}}\exp(\mathbf{q}_{i}^{\top}\mathbf{k}_{t}/\sqrt{d})}$
In matrix multiplication form, we express SWA as: $\mathbf{Y}_{swa}=\mathrm{softmax}\left(\mathbf{Q}\mathbf{K}^{\top}/{\sqrt{d}\odot\mathbf{M}_{swa}}\right)\mathbf{V}$ where $\mathbf{M}_{swa}$ is a mask matrix enforcing the sliding window and the meta-token attention pattern.
These special meta-tokens act as compressed representations that all tokens can attend to globally. Although these meta-tokens are not semantically meaningful, they receive high attention weights from other tokens in the sequence, as also observed in previous studies (Lin et al., 2024; Xiao et al., 2023). Our experiments show that these meta-tokens significantly enhance the model’s recall capability. We illustrate the attention pattern of Lizard in Figure 2. This formulation enables $O(L(w+m)d)$ time and space complexity, where $L$ is the sequence length, $w$ is the window size and $m$ is the number of meta-tokens. Importantly, it also enables constant-memory usage during inference by maintaining only a fixed size $w+m$ key-value cache.
Attention Approximation and Training Objective We approximate the full softmax attention output $\mathbf{Y}_{attn}$ by combining the outputs of gated linear attention and sliding window attention:
$\hat{\mathbf{Y}}_{attn}\approx\hat{\mathbf{Y}}_{gate}+\alpha\cdot\hat{\mathbf{Y}}_{SWA}$
The parameters $\phi_{q}$, $\phi_{k}$, $W_{\gamma}$, and $\alpha$ are optimized via the following mean squared error loss, minimizing the discrepancy between the true and approximated attention outputs:
$\mathcal{L}_{\text{MSE}}(\phi_{q},\phi_{k},W_{\gamma},\alpha)=\frac{1}{N}\sum_{l=1}^{N}\left\|\mathbf{Y}_{attn}^{l}-\left(\hat{\mathbf{Y}}_{gate}^{l}+\alpha\cdot\hat{\mathbf{Y}}_{SWA}^{l}\right)\right\|_{F}^{2}$
where $N$ is the number of attention layers. Overall, Lizard achieves an $O\bigl{(}L(w+m)d+Ld^{2}\bigr{)}$ time and space complexity. For generation, Lizard requires only $O((w+m)^{2}d+d^{2})$ time and space for every token.
3.2 Second Stage: Aligning Subquadratic Attention with Language Modeling
While the first stage focuses on minimizing the approximation error between the hybrid subquadratic attention and full softmax attention, it does not directly optimize for the final language modeling performance. To bridge the gap between architectural approximation and practical performance, we fine-tune the model on autoregressive language modeling objectives to align the approximated attention mechanism with downstream performance. The loss function used during this stage is the standard autoregressive language modeling objective: $\mathcal{L}_{\text{LM}}(\theta)=-\sum_{i=1}^{L}\log P(x_{i}\mid x_{1:i})$
Table 1: Performance comparison of Lizard and existing 7B-size subquadratic LLMs. Linearized models are categorized as Bounded (limited to context length) or Unbounded (capable of extrapolating to longer sequences).
Model Training PiQA ARC-e ARC-c Hella. Wino. MMLU Avg. Avg.
Tokens (B) acc acc acc_norm acc_norm acc (5-shot) (no MMLU)
Transformer
Gemma-7B 6000 81.9 81.1 53.2 80.7 73.7 62.9 74.1 72.3
Mistral-7B 8000* 82.1 80.9 53.8 81.0 74.0 62.4 74.4 72.4
LLaMA-3-8B 15000 79.9 80.1 53.3 79.1 73.1 66.6 73.1 72.0
Subquadratic
Mamba-7B 1200 81.0 77.5 46.7 77.9 71.8 33.3 71.0 64.7
RWKV-6-v2.1-7B 1420 78.7 76.8 46.3 75.1 70.0 69.4 69.4
TransNormerLLM-7B 1400 80.1 75.4 44.4 75.2 66.1 43.1 68.2 64.1
Hawk-7B 300 80.0 74.4 45.9 77.6 69.9 35.0 69.6 63.8
Griffin-7B 300 81.0 75.4 47.9 78.6 72.6 39.3 71.1 65.8
Linearized (Bounded)
Mistral-7B-SUPRA 100 80.4 75.9 45.8 77.1 70.3 34.2 69.9 64.0
Mistral-7B-LoLCATs 0.04 81.5 81.7 54.9 80.7 74.0 51.4 74.5 70.7
LLaMA-3-8B-LoLCATs 0.04 80.9 81.7 54.9 79.7 74.1 52.8 74.2 70.7
Linearized (Unbounded)
Mamba2-LLaMA-3-8B 20 76.8 74.1 48.0 70.8 58.6 43.2 65.6 61.9
Mistral-7B-Lizard (Ours) 0.04 81.8 83.2 55.8 79.8 72.0 60.8 74.5 72.2
LLaMA-3-8B-Lizard (Ours) 0.04 82 83.5 56.7 79.3 71.7 61.2 74.6 72.4
4 Experimental Study
In this section, we present our experimental results, focusing on three key aspects:
  1. 1.
    Language Modeling Benchmarks: We evaluate Lizard on standard language modeling datasets and compare its performance against existing subquadratic alternatives and linearizations. Our results indicate that Lizard matches the average performance of the teacher model and significantly outperforms other baselines by a large margin.
  2. 2.
    Language Modeling Recall: To evaluate memory retrieval capabilities, we conduct experiments on the Needle-in-a-Haystack test, which assesses the model’s ability to retrieve specific information over long sequences. Lizard significantly outperforms existing linearized LLMs on associative recall tasks. Furthermore, by retaining a small number of full attention layers, Lizard closely approximates the teacher model’s performance in the 5-shot MMLU benchmark.
  3. 3.
    Generation Efficiency: We compare the generation throughput of Lizard with the teacher model equipped with FlashAttention-2. While the teacher model quickly runs out of memory at a sequence length of 32K, Lizard maintains constant memory usage and throughput, enabling efficient generation with infinite context.
  4. 4.
    Architectural and Ablation Analysis: We conduct a detailed analysis of Lizard’s design choices, including the structure of the gating module and the contribution of each component. These studies help identify the most effective configurations for performance and efficiency.
Experimental Setup: We conduct our experiments using two widely used Transformer-based LLMs: Mistral-7B (Jiang et al., ) and Llama-3-8B (Grattafiori et al., 2024) as teacher models. For training, we utilize a curated subset of 50K high-quality examples cleaned Alpaca dataset  [^2] https://huggingface.co/datasets/yahma/alpaca-cleaned (Taori et al., 2023). By default, we use a scalar gating structure, where $\Gamma_{i}=\gamma_{i}\mathbf{1}_{d}^{\top},\gamma_{i}=\sigma(W_{\gamma}\mathbf{x}_{i})$, with $\gamma_{i}\in\mathbb{R}$ and $W_{\gamma}\in\mathbb{R}^{d\times 1}$. We also explore various gating module designs, which are discussed in Section 4.4.
In the First Stage, we train the feature maps $\phi_{q}$, $\phi_{k}$, and the gating parameter matrix $W_{\gamma}$ jointly to approximate full softmax attention and RoPE patterns. For the sliding window attention module, we use a small window size of $w=128$ and $m=4$ meta tokens. In the Second Stage, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2022) for parameter-efficient fine-tuning. Specifically, LoRA is applied to the projection matrices $W_{Q}$, $W_{K}$, and $W_{V}$, with a default rank $r=8$ and scaling factor $\alpha=16$. Both stages are trained for 2 epochs, corresponding to 20M tokens per stage. We use the AdamW optimizer and a cosine learning rate schedule with a 10% linear warmup. Training is performed using Fully Sharded Data Parallelism (FSDP-2) (Zhao et al., 2023) across 8$\times$×A100 80GB GPUs. The peak learning rate is set to $1\times 10^{-3}$ for the first stage and $5\times 10^{-4}$ for the second stage. We adopt a global batch size of 8, with each example having a maximum sequence length of 2048 tokens.
4.1 Language Modeling Benchmarks
Table 2: Comparison of hybrid softmax models on language understanding benchmarks.
Model Training PiQA ARC-e ARC-c Hella. Wino. MMLU Avg. Avg.
Tokens (B) acc acc acc_norm acc_norm acc (5-shot) (no MMLU)
LLaMA-3-8B 15000 79.9 80.1 53.3 73.1 79.1 66.6 73.1 72.0
Hybrid Softmax
StripedHyena-Nous-7B 78.8 77.2 40.0 66.4 76.4 26.0 67.8 60.8
Zamba-7B 1000 81.4 74.5 46.6 76.4 80.2 57.7 71.8 69.5
Linearized (Keep 50% Full Attn.)
Mamba2-LLaMA-3 20 81.5 78.8 58.2 71.5 79.5 56.7 73.9 71.0
LLaMA-3-8B-Lizard (Ours) 0.04 82.2 83.1 55.9 73.6 81.4 65.1 75.2 73.5
We evaluate Lizard on six popular language understanding benchmarks from the LM Evaluation Harness (LM Eval)  [^3] https://github.com/EleutherAI/lm-evaluation-harness (Gao et al., 2024), including PiQA (Bisk et al., 2020), ARC-easy (ARC-e) and ARC-challenge (ARC-c) (Clark et al., 2018), HellaSwag (Hella.) (Zellers et al., 2019), WinoGrande (Wino.) (Sakaguchi et al., 2021), and MMLU (Hendrycks et al., 2020). Notably, Lizard is able to closely recover the performance of the teacher model and achieves near-lossless accuracy across tasks in average, demonstrating that it preserves the original model’s language understanding capabilities.
We compare Lizard against two groups of baselines. The first group, presented in Table 1, consists of subquadratic LLMs, including models pre-trained from scratch with linear or subquadratic attention mechanisms, such as Mamba (Gu and Dao, 2023), RWKV-6 (Peng et al., 2023a), TransNormerLLM-7B (Qin et al., ), Hawk and Griffin (De et al., ), as well as linearized variants such as SUPRA (Mercat et al., 2024), LoLCATs (Zhang et al., 2024), and Mamba2LLaMA (Wang et al., 2024). Lizard consistently outperforms prior approaches, particularly on the 5-shot MMLU benchmark, where it achieves an 18% improvement over previous methods with similar extrapolation capabilities. Compared to LoLCATs, which does not generalize beyond training context, Lizard scores 8 points higher on the 5-shot MMLU.
The second group shown in Table 2, Hybrid Softmax Architectures includes models that combine full softmax with subquadratic attention layers. We compare with models such as StripedHyena-Nous-7B (Poli et al., 2023) and Zamba-7B (Glorioso et al., 2024). Following the same configuration of Mamba2-LLaMA-3-8B (Wang et al., 2024), which retains 50% softmax layers. On 5-shot MMLU, Lizard score $65.1$, closely matching the 66.1 score of the original LLaMA-3-8B teacher model, while outperforming all hybrid baselines.
4.2 Recall Evaluations
Figure 3: Example from the synthetic passkey retrieval dataset.
Figure 4: Needle-in-a-Haystack evaluation. Each cell shows retrieval accuracy by sequence length (X-axis) and target distance (Y-axis). Green indicates success; red indicates failure. The white dashed line marks the max training length.
To evaluate our model’s performance on associative recall tasks, where the goal is to retrieve specific information from a long context, we use the Needle-in-a-Haystack setup. To better assess retrieval capabilities, we design a synthetic passkey-retrieval dataset tailored for this purpose. As illustrated in Figure 3, each input sequence contains five randomly generated passkeys, each of length 5-8 tokens, inserted at random positions within a long sequence. At test time, the model is prompted to retrieve one selected passkey from the five embedded within the sequence. We generate 10,000 synthetic examples, train the model on sequences of length 2048, and evaluate its performance on longer sequences ranging from 2048 to 8192 tokens to assess its generalization and recall capabilities in long-context settings.
We evaluate Lizard and compare its performance against the teacher model and LoLCATs (Zhang et al., 2024), a recent state-of-the-art linearization method. Figure  4 reports the results of all three models on the associative recall test set. We find that Lizard significantly outperforms LoLCATs in both associative recall accuracy and length generalization. Notably, Lizard is able to perfectly retrieve the correct passkey across various context lengths, while LoLCATs fails when the sequence length exceeds the training window. This highlights the strength of the gated recurrent structures, which effectively compresses global context and does not rely solely on the expressiveness of local sliding window attention.
Figure 5: Throughput and memory comparison.
4.3 Generation Efficiency
We assess the efficiency of Lizard by comparing its throughput and memory usage to that of the teacher model across input sequence lengths from 1K to 32K, using a batch size of 16. As shown in Figure 5, the teacher model with FlashAttention-2 (Dao, 2023) encounters out-of-memory (OOM) issues at longer sequence lengths. In contrast, Lizard maintains constant memory consumption and stable throughput throughout. All experiments were conducted on an NVIDIA A100 80GB GPU.
Table 3: Performance comparison of different gating designs and their parameterizations.
Model Gating Parameterization Learnable Parameters MMLU 5-shot
Lizard (Ours) $\Gamma_{i}=\gamma_{i}\mathbf{1}_{d}^{\top},\;\gamma_{i}=\sigma(W_{\gamma}\mathbf{x}_{i})$ $W_{\gamma}\in\mathbb{R}^{d\times 1}$ 61.2
Mamba-2 (Dao and Gu, 2024) $\Gamma_{i}=\gamma_{i}\mathbf{1}_{d}^{\top},\;\gamma_{i}=\exp\left(-\operatorname{softplus}(\mathbf{x}_{i}W_{\gamma})\cdot\exp(a)\right)$ $W_{\gamma}\in\mathbb{R}^{d\times 1},\;a\in\mathbb{R}$ 57.6
GLA (Yang et al., 2023) $\Gamma_{i}=\sigma\left(\mathbf{x}_{i}W_{\gamma_{1}}W_{\gamma_{2}}\right)$ $W_{\gamma_{1}}\in\mathbb{R}^{d\times 16},\;W_{\gamma_{2}}\in\mathbb{R}^{16\times d}$ 53.5
1D-Pooling $\Gamma_{i}=\sigma(Pooling(\textbf{k}_{t}))$ N/A 44.1
4.4 Ablation and Architectural Analysis
Gated Structures Design Table 3 presents a comparison of different gating designs and parameterizations based on recent architectural advances. We experiment with multiple formulations, ranging from minimal scalar gates to more expressive multi-layer projections. Our results show that the Lizard parameterization achieves the highest performance on the 5-shot MMLU benchmark. While complex gated recurrent structures offer greater modeling flexibility, we observe that their effectiveness is limited by the need to initialize these modules from scratch. Heavier parameterizations can lead to overfitting or instability during fine-tuning, ultimately degrading performance. In contrast, lightweight designs with minimal additional parameters are easier to train and generalize better, resulting in stronger overall performance. Additionally, we evaluate a pooling-based variant where the gating values are derived from 1D pooling over key vectors, eliminating the need for any learnable parameters. However, this configuration results in a significant drop in performance. This suggests that having a learnable gating mechanism, even with minimal parameters, is crucial for capturing meaningful temporal patterns and maintaining strong performance on downstream tasks.
Table 4: 5-shot MMLU performance with varying window and meta token sizes
$w=32$ $w=64$ $w=128$ $w=256$
$m=2$ 51.3 54.1 58.6 42.2
$m=4$ 52.4 57.6 61.2 44.6
$m=6$ 52.4 57.9 60.8 43.8
Effect of window and meta memory size Table 4 presents an ablation study evaluating the impact of varying the local attention window size ($w$) and the number of meta tokens ($m$) on 5-shot MMLU performance. While increasing the window size can improve performance, it does not guarantee consistent gains. For instance, performance peaks at $w=128$ for $m=4$, but drops notably at $w=256$. This indicates that simply enlarging the local attention span is not always beneficial.
5 Conclusion
We present Lizard, a linearization framework that transforms pretrained Transformers into scalable, subquadratic architectures. Our method achieves near-parity with the teacher model on standard language modeling benchmarks, including 5-shot MMLU, while enabling constant-memory inference and strong length generalization. Through extensive experiments and ablations, we demonstrate that lightweight gating designs and hybrid attention improve both performance and efficiency, making Lizard a practical solution for long-context language modeling.
Appendix A Limitations
Despite the promising performance and efficiency gains demonstrated by Lizard, our approach has two key limitations. First, Lizard still relies on a strong pretrained backbone to achieve high quality. As with many recent distillation-based or hybrid architectures, the success of our method depends heavily on the expressiveness and generalization capacity of the teacher model. Without access to a high-quality pretrained model (e.g., Llama-3-8B), the performance of Lizard may degrade significantly, especially on complex reasoning and multilingual tasks. Second, Lizard inherits the inherent tradeoffs present in linear attention mechanisms. While our design enables constant-time and constant-memory inference with infinite context length, it still exhibits a recall-memory tradeoff. That is, models with fixed-size state representations, such as our gated linear attention, may underperform in recall-intensive tasks compared to full attention models, which maintain a growing key-value cache. This aligns with recent findings that efficient alternatives to attention often struggle to retain long-range information critical for grounding generations in earlier context. As a result, while Lizard expands the throughput-recall Pareto frontier, it does not eliminate the tradeoff entirely.
Appendix B Inference Efficiency
Figure 6: Inference speed comparison between GLA and Lizard kernel..
Hardware-aware GLA in Lizard. We benchmark the Lizard kernel under BF16 precision with batch size $B=16$, sequence length $S=2048$, number of heads $H=32$, and head dimension $D_{\text{head}}=128$. As shown in Figure 6, our hardware-aware implementation of GLA achieves 3.25 ms per forward pass, representing a 32% reduction in inference time compared to the original Gated Linear Attention  [^4] https://github.com/fla-org/flash-linear-attention kernel (4.30 ms). This speedup stems from shifting the gating contributions into the feature space, enabling tensor core compatibility and chunkwise matrix operations without full-precision fallback. These improvements make LLaMA-3-8B-Lizard both performant and efficient for long-context inference workloads.
Appendix C Ablation Study
Impact of Architectural Components We conduct an ablation study to evaluate the contribution of each module in Lizard. As shown in Table 5, removing the sliding window attention (SWA), gated module, or attention approximation significantly degrades performance on both the 5-shot MMLU benchmark and the average across other language understanding tasks. Interestingly, replacing LoRA with full fine-tuning results in similar performance, indicating the effectiveness of parameter-efficient adaptation.
Table 5: Ablation results on MMLU (5-shot) and average performance across other tasks. We report the impact of removing individual components from Lizard.
Model MMLU (5-shot) $\uparrow$ Avg. (no MMLU) $\uparrow$
LLaMA-3-8B-Lizard 61.2 74.6
- SWA 39.7 72.9
- Gated Module 42.2 70.1
- Full FT (no LoRA) 61.4 74.6
- w/o Attention Approximation 50.8 71.2
Effect of LoRA Rank We study how varying the LoRA rank impacts model performance across downstream tasks. As shown in Table 6, we find that a LoRA rank of 8 is sufficient to achieve strong performance, matching or even surpassing full fine-tuning on both 5-shot MMLU and the average across other tasks.
Table 6: Effect of LoRA Rank.
LoRA Rank MMLU (5-shot) $\uparrow$ Avg. (no MMLU) $\uparrow$
4 59.7 74.1
8 61.2 74.6
16 60.6 73.3
32 61 74.5
64 59.2 74
Appendix D Experimental Details
Table 7: Hyperparameters for the experiments.
Resources 8xA100 80GB
Distributed Setup Fully Sharded Data Parallel (FSDP-2)
Model
Precision bfloat16
Sequence length 2048
Hedgehog Feature Dimension 128
Hedgehog Feature Activation Softmax
Optimizer and LR Schedule
Optimizer AdamW
$(\beta_{1},\beta_{2}),\epsilon$ $(0.9,0.99)$, 1.0e-8
Learning Rate min ratio 0.1
Global batch size 8
Micro batch size 1
Gradient Clipping 1.0
Learning rate Rchedule Cosine Annealing LR
Stage 1: Attention Approximation
Number of tokens 20M
Peak Learning Rate $1.0\times 10^{-3}$
Stage 2: Fine-tuning
Number of tokens 20M
Peak Learning Rate $5.0\times 10^{-4}$
LoRA rank and alpha $r=8,\alpha=16$
LoRA dropout 0.0
LoRA projections $W_{q},W_{k},W_{v}$
Hyperparameters All model and training hyperparameters are summarized in Table 7. For the learning rate, we performed an initial sweep over {1e-2, 5e-3, 1e-3, 5e-4, 1e-4}. We did not tune the batch size. For the other designs, we adopted the default values used by prior work [Zhang et al., 2024].
Appendix E Evaluation on small-size LLMs
Table 8: Evaluation results of small-size LLMs and their variants across multiple benchmarks. Lizard consistently outperforms LoLCATs and closely matches the performance of the teacher models.
Model PiQA ARC-e ARC-c Hella. Wino. MMLU Avg. Avg.
acc acc acc_norm acc_norm acc (5-shot) (no MMLU)
Llama-3.2-1B 74.1 65.4 36.4 63.8 60.0 31.0 59.9 55.1
$\rightarrow$ LoLCATS 74.6 63.0 35.1 63.7 61.5 27.3 59.6 54.2
$\rightarrow$ Lizard 74.8 65.6 36.5 64.1 59.7 29.8 60.1 55.1
Llama-3.2-3B 76.4 74.7 46.0 73.6 69.9 56.2 68.1 66.1
$\rightarrow$ Lizard 76.8 75.2 45.2 74.3 69.3 53.4 68.2 65.7
Appendix F Sample Generations
Prompt
Why did we linearize the chungus among us?
LoLCATs Response [Zhang et al., 2024] Linearizing the chungus among us is a process of organizing and arranging the information in a logical and coherent manner. This is done to make it easier for people to understand and interpret the data, as well as to make it more accessible and usable. By linearizing the data, we are able to extract meaningful insights and patterns from the information, which can then be used to make informed decisions and take action.
Lizard Response I’m sorry, but I don’t have any information about linearizing the "Chungus Among Us." Could you please provide more context or clarify what you mean by "linearize the chungus among us"?
Softmax Attention I’m sorry, but I cannot respond to this question as I do not have the necessary information to provide a meaningful response. Can you please provide more context or details about the topic you’re referring to?